Can you please help me by showing me how to simplify this.
((3x^a y^b)^3/ (-3x^a y^b)^2 )^2
I am not sure if I am writing this properly. The problem is written as a fraction and the entire problem is squared.

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

[(3^ay^b)^3]=27(x^3a)(y^3b)
[(-3x^ay^b)^2]=9(x^2a)(y^2b)
[27(x^3a)(y^3b)/9(x^2a)(y^2b)]^2
[3(x^a)(y^b)]^2
9(x^2a)(y^2b) answer.

Answer by stanbon(57239)

((3x^a y^b)^3/ (-3x^a y^b)^2 )^2
------------------------------
[(27x^(3a)y^(3b)) / (9x^(2a)y^(2b)]^2
= [3x^(a)y^(b)]^2
= 9x^(2a)y^(2b)
=============
Cheers,
Stan H>